OpenStudy (anonymous):

5x-1)/2+(-3x-1)/4= 9/2

OpenStudy (anonymous):

\[ \frac{5x-1}{2}+\frac{-3x-1}{4}=\frac{9}{2}\]

OpenStudy (anonymous):

I'm going to assume you left off an opening parenthesis, so we have: (5x-1)/2+(-3x-1)/4= 9/2 We're going to multiply both sides by 4, so we can get rid of the fractions. 2(5x-1) + (-3x-1) = 2*9 Now get rid of the parenthesis, and simplify things. 10x-2 -3x -1 = 18 Let's get all the numbers on one side and all the x's on the other 10x-2-3x-1 = 18 10x-3x = 18 +2+1 7x = 21 x = 3

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.